# MotifDisco: Motif Causal Discovery For Time Series Motifs

### v1

@@ Abstract
Many time series, particularly health data streams, can be best understood as a sequence of phenomenon or events, which we call motifs. A time series motif is a short trace segment which may implicitly capture an underlying phenomenon within the time series. Specifically, we focus on glucose traces collected from continuous glucose monitors (CGMs), which inherently contain motifs representing underlying human behaviors such as eating and exercise. The ability to identify and quantify causal relationships amongst motifs can provide a mechanism to better understand and represent these patterns, useful for improving deep learning and generative models and for advanced technology development (e.g., personalized coaching and artificial insulin delivery systems). However, no previous work has developed causal discovery methods for time series motifs. Therefore, in this paper we develop MotifDisco (motif discovery of causality), a novel causal discovery framework to learn causal relations amongst motifs from time series traces. We formalize a notion of Motif Causality (MC), inspired from Granger Causality and Transfer Entropy, and develop a Graph Neural Network-based framework that learns causality between motifs by solving an unsupervised link prediction problem. We also integrate MC with three model use cases of forecasting, anomaly detection and clustering, to showcase the use of MC as a building block for other downstream tasks. Finally, we evaluate our framework and find that Motif Causality provides a significant performance improvement in all use cases.
@@ 1 Introduction
Many time series can be best understood as sequences of phenomenon or events. This is extremely common in many health data streams where traces are guided by underlying human physiology or behaviors. We call these events in the traces motifs. A time series motif is a short trace segment which may implicitly capture an underlying behavior or phenomenon within the time series. To contextualize our discussion, and as our main running example, we focus on glucose traces collected from continuous glucose monitors (CGMs) for diabetes. Glucose traces inherently contain motifs which represent underlying human behaviors. For instance, a motif capturing a peak in glucose may correspond to an individual eating; a motif capturing a drop in glucose may correspond to an individual exercising. Figure 1 shows some real sample glucose traces and motifs.
Causal discovery is the process by which causal relations are found amongst observational data (Niu et al., 2024). The ability to discover and quantify causality amongst motifs can provide a mechanism to better understand and represent these patterns, useful in a variety of applications. For instance, learning causal relations in glucose motifs may enable better understanding of physiological patterns contributing to advanced technology development (e.g., artificial insulin delivery systems). Moreover, motif causal relationships may be helpful building blocks when used as sub-components in generative and deep learning models to improve their performance for other downstream tasks.
Granger Causality (Granger, 1969) and its nonlinear extension Transfer Entropy (TE) (Schreiber, 2000) are commonly used in time series causal discovery methods, as causal relations are quantified based on one trace’s predictability of another. This notion of causality provides an intuitive, understandable measure to quantify causal relations and is advantageous for time series because it innately incorporates temporality without requiring strong model assumptions. Despite exciting recent developments for time series causal discovery (Gong et al., 2023; Assaad et al., 2022), no previous work has focused on causal discovery for time series motifs. Motif causal discovery is challenging because in many cases (particularly in health) there is no ground truth about what underlying behavior a motif captures. For example, from data alone one cannot conclusively determine what caused a change in glucose (e.g., a glucose rise may be from eating or stress). As a result, unlike in many other causal models, there is no ground truth causal structure amongst motifs that could be used to guide the casual discovery model (i.e., through supervised methods using labeled causal events.)
@@ 2 Related Work
Motifs. Recently, there has been interest in temporal network motifs, which are sets of recurring graph substructures. Previous work has investigated causality within network motifs (Liu et al., 2021; Kovanen et al., 2011); developed network motif causal discovery and graph representation learning frameworks (Chen & Ying, 2024; Chen et al., 2023b; Jin et al., 2022); and developed new use case architectures that utilize network motif causal structures, such as for pharmacovigilance and anomaly detection tasks (Xiao et al., 2024; Kalla et al., 2023; Yu & Gao, 2022; Zhu et al., 2022; Mu et al., 2021). Importantly, the definition of motif used here is different than ours, referring to patterns in graph structures as opposed to patterns in the traces themselves. In other motif application areas, Chinpattanakarn & Amornbunchornvej (2024) solve a different problem and develop a method that, given two time series, infers a set of patterns, also called motifs, that follow each other in the traces. Finally, Lamp et al. (2024) develop a method to generate realistic synthetic time series glucose traces and use a notion of causality amongst motifs to help the model perform well. The focus of this work is not on causality, and the causal learning method is complicated and suffers from scalability issues.
Causal Discovery in Time Series. There are a variety of works on causal discovery for time series (Niu et al., 2024; Gong et al., 2023; Hasan et al., 2023; Assaad et al., 2022; Shojaie & Fox, 2022). In particular, previous methods have developed causal discovery frameworks for multivariate time series that incorporate temporal dynamics and use Granger Causality (Pan et al., 2024; Löwe et al., 2022) or Transfer Entropy (Bonetti et al., 2024; Najafi et al., 2023). Recent work has also incorporated time series causality measures to improve model learning in downstream tasks such as forecasting and anomaly detection (Ansari et al., 2024; Chen et al., 2023a; Febrinanto et al., 2023; Duan et al., 2022; Wu et al., 2021). MotifDisco is the first framework focused on causal discovery for time series motifs, which may broaden current model capabilities for event-based time series.
@@ 3 Formalizing Motif Causality
Motifs & Motif Construction. We first formalize our notion of motifs. A motif is a short segment of the trace which may implicitly capture an underlying behavior within a time series. Real sample glucose traces and motifs are shown in Figure 1. We define a motif, $\mu$, as a short, ordered sequence of values ($v$) of length $\tau$:
@@
We denote a set of $n$ time series traces as $X=[x^{1},\ldots,x^{n}]$. Each time series may be represented as a sequence of motifs: $x^{j}=[\mu_{1}^{i},\mu_{2}^{i},\ldots]$ where each $\mu^{i}_{t}$ gives the motif identifier $i$ at the ordered time step $t$. We also define a motif set $\mathcal{M}$, of $|m|$, which is the complete set of motifs generated from the traces =$\mathcal{M}=\{\mu^{1},\ldots,\mu^{m}\}$. The user may choose how they wish to pull out motifs from the time series based on the end application goal. We assume there is a consistent, conclusive way to pull motifs from the traces, and motif discovery is outside the scope of this work. We refer the interested reader to other works focused on this problem (Chinpattanakarn & Amornbunchornvej, 2024; Schäfer & Leser, 2022; Ye & Keogh, 2009). That being said, three straightforward methods to extract motifs from traces, shown in Figure 2, include chopping the traces into size $\tau$ chunks 2(a), using a sliding window of size $\tau$ to extract motifs 2(b), and using signal processing techniques such as Discrete Fourier Transforms (DFT) to automatically extract motifs 2(c).
Granger Causality & Transfer Entropy. Granger Causality is a common method to characterize causal relations amongst time series (Granger, 1969). Different from other causal methods, Granger defines causal relations in terms of predictability. Under Granger Causality, given two time series $x$ and $y$, $x$ causes $y$ if past information about $x$ is more predictive than past information about $y$ only. Transfer Entropy (TE), sometimes also called Causation Entropy, is a nonlinear extension of Granger Causality (Schreiber, 2000; Barnett et al., 2009). Using information theoretic measures, TE measures the amount of uncertainty that is reduced in future states of time series $y$ as a result of knowing the past states of time series $x$. Given two time series, $x^{i}$ and $y^{i}$, the TE from $x^{i}$ to $y^{i}$ is:
@@
Traditional TE is under the important assumption that the effect is influenced by the cause under a fixed, constant time delay. However, this assumption does not hold for many real world time series applications, particularly in health streams, where data may be affected by past events at varying lengths of time. As such, an extension of TE that allows for different time delays has been developed, denoted here as Variable-lag Transfer Entropy (VTE) (Amornbunchornvej et al., 2021). The VTE from $x^{i}$ to $y^{i}$ is defined as:
@@
Defining Motif Causality. Our definition of motif causality is inspired from various Transfer Entropy and Causation Entropy threads (Equation 3, Irribarra et al. (2024); Gong et al. (2023); Amornbunchornvej et al. (2021); Assaad et al. (2021); Sun et al. (2015)). The Motif Causality (MC) from motif $\mu^{i}$ to motif $\mu^{j}$ conditioned on the set of motifs $\mathcal{K}$ is defined as:
@@
where $\mathcal{K}\subset\mathcal{M}$, $H(\cdot|\cdot)$ is a conditional entropy function, $f$ and $g$ are lag constants and $\Delta_{t}$ is a variable length lag amount. Essentially, this provides a measure of information gain by determining how much uncertainty is reduced for $\mu^{j}$ by observing past occurrences of $\mu^{i}$ compared to the “status quo”, the set of the rest of the motifs $\mathcal{K}$. The $MC$ value will be between 0.0 and 1.0. Higher values indicate stronger causality (i.e., more uncertainty about the future is reduced for $\mu^{j}$ given $\mu^{i}$).
Conditional Entropy Function. To implement the conditional entropy function, $H(\cdot|\cdot)$, there are many different types of entropy which the user can choose based on their end goal or application. For our purposes, we elucidate two common ones: Shannon entropy (Shannon, 1948) and Rényi entropy (Jizba et al., 2012; 2022). Shannon entropy is defined as:
@@
where $\alpha$ is a weight parameter, $\alpha>0$. When $\alpha\rightarrow 1$ Rényi entropy converges to Shannon entropy. We note that there are many other types of entropy functions that could be used and might be relevant, such as Wavelet (Rosso et al., 2001) or Permutation Entropy (Bandt & Pompe, 2002).
@@ 4.1 Preliminaries
Figure 4: Model Architecture: a GNN consisting of stacks of GraphSAGE, ReLU, Dropout (“X” boxes) and Linear layers learns a node embedding. The LinkPredictor consists of Linear, ReLU and Dropout layers followed by the “MC” motif causality layer and a Sigmoid layer ($\sigma$ box). The LinkPredictor takes in the node embedding and outputs the predicted edges and their weights.
@@ 4.2 Model Architecture
GNN. The Graph Neural Network takes in a starting motif graph $G$, and outputs a learned node embedding. The GNN is structured using GraphSage convolutional layers. GraphSAGE (standing for Sample and Aggregate) (Hamilton et al., 2017) learns low dimensional vector representations of nodes. The core intuition of GraphSAGE is that a node is known by the company it keeps (i.e., its neighbors). The algorithm works by iterating over a sample of the node’s neighboring nodes and “aggregating” their embeddings in order to determine the current node’s embedding. Importantly, GraphSAGE is inductive, meaning it can generalize to unseen nodes. GraphSage use both node features and topological structure (i.e., the graph structure) of each node’s neighbourhood simultaneously to efficiently generate representations for new nodes without requiring model retraining. To do this, the algorithm relies on its aggregation function, also known as the message-passing process, which learns how to aggregate node features based on encoding information about a node’s local neighborhood. Therefore, when given new node data, the function uses the local neighborhood of the node to aggregate the features appropriately, and learn the embedded feature representation (as opposed to needing to learn a unique embedding for every individual node). As shown in Figure 4, in our framework the GNN consists of sequential stacks of GraphSAGE convolutional, ReLU and dropout layers (represented by the “X” boxes in the figure) followed by the message-passing layers– stacks of linear and dropout layers. The motif time series values are the node data used to learn the embeddings. We instantiate the aggregation function using mean aggregation for simplicity.
(a) Large Uncertainty Reduction
(b) Small Uncertainty Reduction
Figure 5: Depiction of Conditional Entropy $H(\mu)$ computation using the histogram method: when computing MC for $\mu^{i}$, 5(a) $\mu^{j}$ results in a large uncertainty reduction; 5(b) $\mu^{z}$ has a small reduction.
@@
Link Predictor. The LinkPredictor network takes in learned node embeddings outputted from the GNN and a list of edges to predict, and returns the probability of each edge and the motif causality values for each edge. At its core, the LinkPredictor learns a function to predict the probability of an edge between two nodes. To do this, it computes the MC between the nodes using Equation 4, and a probability score, represented by the element-wise dot product of the two embedded node vectors. In terms of architecture, the LinkPredictor is implemented via stacks of linear, ReLU and dropout layers followed by Motif Causality ($MC$) and sigmoid layers to learn the edge probability function. The network balances evaluating the product of the embedded node vectors and the motif causal values between the motifs themselves, allowing the network to learn how to optimize edge addition/deletion in the graph guided by the underlying causality. (More details about this in subsection 4.3). In this way, the LinkPredictor learns to predict edges that have high motif causality with high probability.
To implement the MC computation, we adapt existing Transfer Entropy libraries (Behrendt et al., 2019) and compute the conditional entropy function (i.e., $H(\mu)$, Shannon’s or Rényi entropy) using the histogram method. A simplified depiction is shown in Figure 5. Essentially, motif time series values are binned into a histogram. Distributions between motifs can be compared to determine how much uncertainty about future predictions of the motif is reduced in the distribution. For example, in 5(a) when computing MC for $\mu^{i}$, $\mu^{j}$ covers a larger distribution, resulting in a large reduction in uncertainty and higher motif causality, whereas 5(b) $\mu^{z}$ covers hardly any new distribution compared to the status quo ($H(\mu^{i}|\mathcal{K})$), resulting in a small reduction in uncertainty and low causality for $\mu^{z}$.
4.3 Model Training

### v2

@@ Abstract
Many time series, particularly health data streams, can be best understood as a sequence of phenomenon or events, which we call motifs. A time series motif is a short trace segment which may implicitly capture an underlying phenomenon within the time series. Specifically, we focus on glucose traces collected from continuous glucose monitors (CGMs), which inherently contain motifs representing underlying human behaviors such as eating and exercise. The ability to identify and quantify causal relationships amongst motifs can provide a mechanism to better understand and represent these patterns, useful for improving deep learning and generative models and for advanced technology development (e.g., personalized coaching and artificial insulin delivery systems). However, no previous work has developed causal discovery methods for time series motifs. Therefore, in this paper we develop MotifDisco (motif discovery of causality), a novel causal discovery framework to learn causal relations amongst motifs from time series traces. We formalize a notion of Motif Causality (MC), inspired from Granger Causality and Transfer Entropy, and develop a Graph Neural Network-based framework that learns causality between motifs by solving an unsupervised link prediction problem. We integrate MC with three model use cases of forecasting, anomaly detection and clustering, to showcase the use of MC as a building block for downstream tasks. Finally, we evaluate our framework on different health data streams and find that Motif Causality provides a significant performance improvement in all use cases.
@@ 1 Introduction
Many time series can be best understood as sequences of phenomenon or events. This is extremely common in many health data streams where traces are guided by underlying human physiology or behaviors. We call these events in the traces motifs. A time series motif is a short trace segment which may implicitly capture an underlying behavior or phenomenon within the time series. To contextualize our discussion, and as our main running example, we focus on glucose traces collected from continuous glucose monitors (CGMs) for diabetes. Glucose traces inherently contain motifs which represent underlying human behaviors. For instance, a motif capturing a peak in glucose may correspond to an individual eating; a motif capturing a drop in glucose may correspond to an individual exercising. Figure 1 shows real glucose traces and motifs (other examples in Appx. A.1).
Causal discovery is the process by which causal relations are found amongst observational data [1]. The ability to discover and quantify causality amongst motifs can provide a mechanism to better understand and represent these patterns, useful in a variety of applications. For instance, learning causal relations in glucose motifs may enable better understanding of physiological patterns contributing to advanced technology development (e.g., artificial insulin delivery systems). Moreover, motif causal relationships may be helpful building blocks when used as sub-components in generative and deep learning models to improve their performance for other downstream tasks.
Granger Causality [2] and its nonlinear extension Transfer Entropy (TE) [3] are commonly used in time series causal discovery methods, as causal relations are quantified based on one trace’s predictability of another. This notion of causality provides an intuitive, understandable measure to quantify causal relations and is advantageous for time series because it innately incorporates temporality without requiring strong model assumptions. Despite exciting recent developments for time series causal discovery [4, 5], no previous work has focused on causal discovery amongst time series motifs. Motif causal discovery is challenging because in many cases (particularly in health) there is no ground truth about what underlying behavior a motif captures. For example, from data alone one cannot conclusively determine what caused a change in glucose (e.g., a glucose rise may be from eating or stress). As a result, unlike in many other causal models, there is no ground truth causal structure amongst motifs that could be used to guide the casual discovery model (i.e., through supervised methods using labeled causal events.)
@@ 2 Related Work
Motifs. Recently, there has been interest in temporal network motifs, which are sets of recurring graph substructures. Previous work has investigated network motif causality [6, 7] and developed network motif causal discovery frameworks [8, 9, 10]. Importantly, the definition of motif used here is different than ours, referring to patterns in graph structures as opposed to patterns in the traces themselves. In other motif application areas, Chinpattanakarn and Amornbunchornvej [11] solve a different problem and develop a method to infer a set of patterns, also called motifs, that follow each other in the traces. Finally, Lamp et al. [12] develop a method to generate synthetic time series glucose traces and use a notion of causality amongst motifs to help the model perform well. The focus of this work is not on causality, and the causal learning method is complicated and suffers from scalability issues.
Causal Discovery in Time Series. There are a variety of works on causal discovery for time series [1, 4, 13, 5, 14]. In particular, previous methods have developed causal discovery frameworks for multivariate time series that incorporate temporal dynamics and use Granger Causality [15, 16, 17] or Transfer Entropy [18, 19]. Recent work has also incorporated time series causality measures to improve model learning in downstream tasks such as forecasting and anomaly detection [20, 21, 22, 23, 24]. Previous methods for time series causal discovery cannot be directly applied to motifs because they formulate causality using multivariate statistical properties (e.g., variable-based correlation) or temporal statistical measures repeated across time series lags, which do not hold for short, univariate time series motifs that do not contain repeated lagged patterns; or use labels or known underlying causal structures, which are not available for our traces and many similar event-based data streams. MotifDisco is the first framework for causal discovery in time series motifs, which may broaden current model capabilities for event-based time series.
@@ 3 Formalizing Motif Causality
Motifs & Motif Construction. We first formalize our notion of motifs. A motif is a short segment of the trace which may implicitly capture an underlying behavior within a time series. Real sample traces and motifs are shown in Figure 1 and Appx. A.1. We define a motif, $\mu$, as a short, ordered sequence of values ($v$) of length $\tau$:
@@
We denote a set of $n$ time series traces as $X=[x^{1},\ldots,x^{n}]$. Each time series may be represented as a sequence of motifs: $x^{j}=[\mu_{1}^{i},\mu_{2}^{i},\ldots]$ where each $\mu^{i}_{t}$ gives the motif identifier $i$ at the ordered time step $t$. We also define a motif set $\mathcal{M}$, of $|m|$, which is the complete set of motifs generated from the traces =$\mathcal{M}=\{\mu^{1},\ldots,\mu^{m}\}$. We assume there is a consistent, conclusive way to pull motifs from the traces; the user may choose which method they use based on the end application goal. Motif discovery is outside the scope of this work since this problem has been widely solved (see [11, 25, 26] and Appx A.2). That being said, three straightforward methods to extract motifs from traces, shown in Figure 2, include chopping the traces into size $\tau$ chunks 2(a), using a sliding window of size $\tau$ to extract motifs 2(b), and using signal processing techniques such as Discrete Fourier Transforms (DFT) to automatically extract motifs 2(c). Our framework is amenable to any motif extraction method.
Granger Causality & Transfer Entropy. Granger Causality is a common method to characterize causal relations amongst time series [2]. Different from other causal methods, Granger defines causal relations in terms of predictability. Under Granger Causality, given two time series $x$ and $y$, $x$ causes $y$ if past information about $x$ is more predictive than past information about $y$ only. Transfer Entropy (TE), sometimes also called Causation Entropy, is a nonlinear extension of Granger Causality [3, 27]. Using information theoretic measures, TE measures the amount of uncertainty that is reduced in future states of time series $y$ as a result of knowing the past states of time series $x$. Given two time series, $x^{i}$ and $y^{i}$, the TE from $x^{i}$ to $y^{i}$ is:
@@
Traditional TE is under the important assumption that the effect is influenced by the cause under a fixed, constant time delay. However, this assumption does not hold for many real world time series applications, particularly in health streams, where data may be affected by past events at varying lengths of time. As such, an extension of TE that allows for different time delays has been developed, denoted here as Variable-lag Transfer Entropy (VTE) [28]. The VTE from $x^{i}$ to $y^{i}$ is defined as:
@@
Defining Motif Causality. Our definition of motif causality is inspired from various Transfer Entropy and Causation Entropy threads (Equation 3, Irribarra et al. [29], Gong et al. [4], Amornbunchornvej et al. [28], Assaad et al. [30], Sun et al. [31]). The Motif Causality (MC) from motif $\mu^{i}$ to motif $\mu^{j}$ conditioned on the set of motifs $\mathcal{K}$ is defined as:
@@
where $\mathcal{K}\subset\mathcal{M}$, $H(\cdot|\cdot)$ is a conditional entropy function, $f$ and $g$ are lag constants and $\Delta_{t}$ is a variable length lag amount. Essentially, this provides a measure of information gain by determining how much uncertainty is reduced for $\mu^{j}$ by observing past occurrences of $\mu^{i}$ compared to the “status quo", the set of the rest of the motifs $\mathcal{K}$. The $MC$ value will be between 0.0 and 1.0. Higher values indicate stronger causality (i.e., more uncertainty about the future is reduced for $\mu^{j}$ given $\mu^{i}$).
Conditional Entropy Function. To implement the conditional entropy function, $H(\cdot|\cdot)$, there are many different types of entropy which the user can choose based on their end goal or application. For our purposes, we elucidate two common ones: Shannon entropy [32] and Rényi entropy [33, 34]. Shannon entropy is defined as:
@@
where $\alpha$ is a weight parameter, $\alpha>0$. When $\alpha\rightarrow 1$ Rényi entropy converges to Shannon entropy. We note that there are many other types of entropy functions that could be used and might be relevant, such as Wavelet [35] or Permutation Entropy [36].
@@ 4.1 Preliminaries
Figure 4: Model Architecture: a GNN consisting of stacks of GraphSAGE, ReLU, Dropout (“X" boxes) and Linear layers learns a node embedding. The LinkPredictor consists of Linear, ReLU and Dropout layers followed by the “MC" motif causality layer and a Sigmoid layer ($\sigma$ box). The LinkPredictor takes in the node embedding and outputs the predicted edges and their weights.
@@ 4.2 Model Architecture
GNN. The Graph Neural Network takes in a starting motif graph $G$, and outputs a learned node embedding. The GNN is structured using GraphSage convolutional layers. GraphSAGE (standing for Sample and Aggregate) [37] learns low dimensional vector representations of nodes. The core intuition of GraphSAGE is that a node is known by the company it keeps (i.e., its neighbors). The algorithm works by iterating over a sample of the node’s neighboring nodes and “aggregating" their embeddings in order to determine the current node’s embedding. Importantly, GraphSAGE is inductive, meaning it can generalize to unseen nodes. GraphSage use both node features and topological structure (i.e., the graph structure) of each node’s neighbourhood simultaneously to efficiently generate representations for new nodes without requiring model retraining. To do this, the algorithm relies on its aggregation function, also known as the message-passing process, which learns how to aggregate node features based on encoding information about a node’s local neighborhood. Therefore, when given new node data, the function uses the local neighborhood of the node to aggregate the features appropriately, and learn the embedded feature representation (as opposed to needing to learn a unique embedding for every individual node). As shown in Figure 4, in our framework the GNN consists of sequential stacks of GraphSAGE convolutional, ReLU and dropout layers (represented by the “X" boxes in the figure) followed by the message-passing layers– stacks of linear and dropout layers. The motif time series values are the node data used to learn the embeddings. We instantiate the aggregation function using mean aggregation for simplicity.
@@
Link Predictor. The LinkPredictor network takes in learned node embeddings outputted from the GNN and a list of edges to predict, and returns the probability of each edge and the motif causality values for each edge. At its core, the LinkPredictor learns a function to predict the probability of an edge between two nodes. To do this, it computes the MC between the nodes using Equation 4, and a probability score, represented by the element-wise dot product of the two embedded node vectors. To implement the MC computation, we adapt existing Transfer Entropy libraries and compute the conditional entropy function, $H(\mu)$, using the histogram method [38]; additional details in Appx. A.2. In terms of architecture, the LinkPredictor is implemented via stacks of linear, ReLU and dropout layers followed by Motif Causality ($MC$) and sigmoid layers to learn the edge probability function. The network balances evaluating the product of the embedded node vectors and the motif causal values between the motifs themselves, allowing the network to learn how to optimize edge addition/deletion in the graph guided by the underlying causality. (More details provided next in subsection 4.3). In this way, the LinkPredictor learns to predict edges that have high motif causality with high probability.
@@ 7       $loss=-\sum_{i=1}^{b}\left(y_{i}\log\hat{y}_{i}+(1-y_{i})\log(1-\hat{y}_{i})\right)$
4.3 Model Training
5 Uses of Learned Motif Causality
To illustrate the use of Motif Causality as a building block for other downstream tasks, in this section we integrate Motif Causality with three example model use cases of Forecasting, Anomaly Detection and Clustering. We evaluate the performance of these use cases later in section 6.